OpenStudy (anonymous):

Part A: The area of a square is (9x2 − 12x + 4) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. (5 points) Part B: The area of a rectangle is (25x2 − 16y2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work. (5 points)

OpenStudy (anonymous):

@Keigh2015

OpenStudy (anonymous):

@Kash_TheSmartGuy

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

Can someone help me

OpenStudy (anonymous):

hi

OpenStudy (mathstudent55):

Part A. Factor 9x2 − 12x + 4 This trinomial follows the pattern \(a^2 - 2ab + b^2 = (a - b)^2\)

OpenStudy (anonymous):

ok im gonna plug it in

OpenStudy (anonymous):

one sec

OpenStudy (mathstudent55):

To factor it, first, what is it whose square is to 9x^2?

OpenStudy (anonymous):

um..

OpenStudy (anonymous):

idk

OpenStudy (mathstudent55):

What do you square to get 9x^2? ___ * ___ = 9x^2 What is the square root of 9?

OpenStudy (mathstudent55):

What number times itself equals 9?

OpenStudy (anonymous):

3

OpenStudy (mathstudent55):

Good. What times itself is x^2?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

nothing

OpenStudy (mathstudent55):

No. 1 * 1 = 1 The answer is x. x * x = x^2 Ok?

OpenStudy (anonymous):

okay

OpenStudy (mathstudent55):

That means (3x)(3x) = 9x^2 The square root of 9x^2 is 3x

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now we deal with the 4. What is the square root of 4?

OpenStudy (anonymous):

2

OpenStudy (mathstudent55):

Good. Since we have the pattern: \(a^2 - 2ab + b^2 = (a - b)^2\) in this case we have: \(9x^2 - 12x + 4 = (3x - 2)^2\)

OpenStudy (mathstudent55):

That means the side of the square is 3x - 2 units

OpenStudy (anonymous):

so thats the length of each side?

OpenStudy (anonymous):

@phi

OpenStudy (mathstudent55):

Yes. Since it's a square, all sides measure 3x - 2

OpenStudy (anonymous):

ooh okay thankyou

OpenStudy (anonymous):

can you help with B?

OpenStudy (mathstudent55):

In part B, you have a rectangle. A rectangle may or may not be a square. We need to factor (25x2 − 16y2)

OpenStudy (anonymous):

(5x+4y)(5x−4y)

OpenStudy (mathstudent55):

Here, the pattern of the binomial is: \(a^2 - b^2 = (a + b)(a - b) \) This is called the difference of two squares.

OpenStudy (mathstudent55):

You are correct. As you can see the sides are not equal, so this rectangle is not a square.

OpenStudy (mathstudent55):

Two opposite sides have length 5x + 4y, and the other two opposite sides have length 5x - 4y